We consider a nuclear reactor of power output P=1000 Megawatt (1000 million watts) electric, functioning with Plutonium. It is fueled, initially, with 1000 kg of Plutonium. The nuclear material in question is made of Plutonium nuclei, each consisting in 94 protons and 239-94=145 neutrons, which is denominated by the symbol 94Pu239. For thermodynamical reasons, only 1/3rd of the nuclear energy in the form of heat produced by the reactor, can be converted into electricity. How much mass deficit Δm should the nuclear fuel of concern delineates, if the reactor is operated through a period of Δt=2 years. You are supposed to use Einstein’s relationship ΔE=c2Δm (where c is the velocity of light in vacuum, i.e. about 300 000 km/s). Note further that the reactor power, by definition, is given by ΔE/Δt.Make sure that you use coherent units.

b) (10 p) How much mass deficit between the reactants (C+O2) and the reaction products (CO2), all weighed at rest, would delineate a coal (C) thermal plant of the same capacity, working under the same thermodynamical regime, through the same period of 2 years of operation, but which needs to burn about 3 millions coal per year?
Answer #1

Given, nuclear reactor power output P = 1000 MW = 1000*10^6 W = 10^9 W
functioning with plutonium
initial mass of plutonium is 1000 kg

now, the nuclear power plant works for 2 years and efficiency is 1/3
hence
Total energy input for 2 years is
Ein = 3P*2 years = 3*10^9*2*365*24*60*60 = 189.216*10^15 J
now from einsteins equation
E = mc^2
189.216*10^15 = m*9*10^16
m = 2.1024 kg
so mass deficit is m = 2.1024 kg

b. For the cola power plant burning 3 million tons of coal per year
m = 3*10^9 kg per year
for 2 years, m = 6*10^9 kg

now, mass of CO2 is 44.01g/mol
mass of C is 12.011 g/mol
mass of O2 is 2*15.9994 g/mol
mass deficit per mole of reaction is 44.01 - 12.011 - 2*15.9994 = 0.0002 g/mol
so, for 6*10^9 kg carbon, number of moles of C is 1.363326516*10^11 mol
total mass deficit is 0.0002*1.3633265167007498295841854124063*10^11 = 27266.5303340 kg
total mass deficit for the same time in cola plant is 27266.5303 kg , which is 12969.240075 times the mass deficit in the nuclear power plant